OpenStudy (anonymous):

will medal for best answer can you answer this [4-(-2)*3}/2

OpenStudy (anonymous):

So just take it in pieces. 4*-2 = -8 -8*3=-24 -24/2=-12 so you answer is -12! Hope this helps!